the volume of a gas at temperature -73°c and pressure 480mm hg is 1000 cm find the volume of the same gas at pressure 760mm of maccury and 127°c


1
Expert's answer
2021-02-25T12:26:40-0500

Let's use the combined gas law:


"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2},""V_2=\dfrac{P_1V_1T_2}{P_2T_1},""V_2=\dfrac{480\ mm\ Hg\cdot\dfrac{133.322\ Pa}{1\ mm\ Hg}\cdot1000\ cm^3\cdot(\dfrac{1\ m}{100\ cm})^3\cdot400.15\ K}{760\ mm\ Hg\cdot\dfrac{133.322\ Pa}{1\ mm\ Hg}\cdot200.15\ K},""V_2=1.263\cdot10^{-3}\ m^3=1263\ cm^3."